Traveling for 3hr into a steady head wind, a plane flies 1650 mi. The pilot determines that flying with the sane wind for 2 hr, he could make a trip of 1300mi. Find the speed of the plane and the speed of the wind.
Answer by checkley77(12569)

1650/3=550 mph into thw wind
1300/2=650
(650-550)/2=100/2=50 mph. is the speed of the wind.
550++50=600 mph. is the speed of the plane in still air.